OpenStudy (anonymous):

What direct current will produce the same amount of thermal energy, in a particular resistor, as an alternating current that has a maximum value of 2.60 A? (Answer: 1.84 A)

OpenStudy (ash2326):

This is the definition of RMS current. RMS or effective value of a current is by definition such that the heating effect is the same for equal values of alternating- or direct-current. Here the ac current's max value is 2.60 A Can you find the rms value of this current ?

OpenStudy (anonymous):

That is I rms with the formula: I rms = Imax/(sqrt(2)) So I rms = Imax/(sqrt(2)) I rms = 2.60 /1.4142 I rms = 1.838 A